OpenStudy (anonymous):

Write the function in the form f(x) = (x−k)q(x) + r for the given value of k. f(x) = x3 + 2x2 − 7x − 9, k = −sqrt(7)

OpenStudy (anonymous):

very ugly. you have to use synthetic division answer is \[ (-2 \sqrt{7}x^2+(2-\sqrt{7}) x-2 \sqrt{7})×(x+\sqrt{7})+5\]

OpenStudy (anonymous):

i used this answer and was given an incorrect mark